A vegetable garden measures 8 ft by 12 ft. By what equal amount must each dimension be increased if the area is to be doubled?
Answer by user_dude2008(1862)

(8+x)(12+x)=2*96


96+20x+x^2=192


x^2+20x-96=0

(x-4)(x+24)=0

x-4=0 or x+24=0

x=4 or x=-24

Answer: 4ft